A 12-V battery is connected across a 100-Ω resistor. How many electrons flow through the wire in 5.0 min?
Answer #1

To calculate the number of electrons flowing through the wire:


Step 1: Calculate the current using Ohm's Law

I=VRI = \frac{V}{R}

Where:

  • V=12VV = 12 \, \text{V}

  • R=100ΩR = 100 \, \Omega

I=12100=0.12AI = \frac{12}{100} = 0.12 \, \text{A}


Step 2: Calculate the total charge that flows in 5.0 minutes

Time:

t=5.0×60=300secondst = 5.0 \times 60 = 300 \, \text{seconds}

Charge:

Q=I×t=0.12×300=36.0CQ = I \times t = 0.12 \times 300 = 36.0 \, \text{C}


Step 3: Calculate the number of electrons

Each electron has a charge of:

e=1.602×1019Ce = 1.602 \times 10^{-19} \, \text{C}Number of electrons=Qe=36.01.602×10192.25×1020electrons\text{Number of electrons} = \frac{Q}{e} = \frac{36.0}{1.602 \times 10^{-19}} \approx 2.25 \times 10^{20} \, \text{electrons}


Final answer: Approximately 2.25×10202.25 \times 10^{20} electrons.


answered by: Monu Kumar Gupta